OpenStudy (anonymous):

If I have the following inequality: -1 <= X <= 1 and Y = 3X How can I change the inequality to be in terms of Y? I wasn't sure how to do this, so I just multiplied -1 times 3, and 1 times 3, and replaced X with Y to get: -3 <= Y <= 3 Now, is this correct? How would I do this for the same inequality: -1 <= X <= 1 but when Y = 3-X? 3-(-1) = 4 3-1 = 2 Replacing X with Y and substituting into the new inequality, I get: 4 <= Y <= 2 which does not make any sense. So what is the method to do this?

OpenStudy (anonymous):

-1<= x <= 1 y = 3-x so x = 3-y -1<= 3-y <= 1 subtract 3 from each -4<=-y <=-2 change signs and reverse inequalities 4>= y >= 2 check y = 3-x, and when x =1 y = 2, when x = -1, y=4 so y is between 2 and 4 when x is between -1 and 1 QED

OpenStudy (anonymous):

@douglaswinslowcooper Thanks. :)

OpenStudy (anonymous):

You are very welcome.